OpenStudy (vshiroky):

For the following question(s): Professor Juarez thinks the students in her statistics class this term are more creative than most students at this university. A previous study found that students at this university had a mean score of 35 on a standard creativity test. Professor Juarez finds that her class scores an average of 40 on this scale, with an estimated population standard deviation of 7. The standard deviation of the distribution of means comes out to 1.63. 4. What is the t score? A) (40  35)/7 = .71 B) (40  35)/1.63 = 3.07 C) (40  35)/72 = 5/49 = .10 D) (40  35)/1.63

OpenStudy (vshiroky):

5. What effect size did Professor Juarez find? A) (40  35)/7 = .71 B) (40  35)/1.63 = 3.07 C) (40  35)/72 = 5/49 = .10 D) (40  35)/1.632 = 5/2.66 = 1.88

OpenStudy (vshiroky):

@jim_thompson5910 Can you double check these? I got 4. B and 5. A

jimthompson5910 (jim_thompson5910):

What's with the squiggle mark over the 40?

OpenStudy (vshiroky):

I don't know it's supposed to be a -

jimthompson5910 (jim_thompson5910):

so like a bar over 40?

OpenStudy (vshiroky):

it's supposed to be (40-35)/7 and so forth

jimthompson5910 (jim_thompson5910):

oh i see

jimthompson5910 (jim_thompson5910):

4 B looks good, but again, I have no idea about the effect size...so idk what to do with 5

OpenStudy (vshiroky):

Ok.. can I show you a few more that I am unsure about?

jimthompson5910 (jim_thompson5910):

go ahead

OpenStudy (vshiroky):

For the following question(s): A school counselor claims that he has developed a technique to reduce prestudying procrastination in students. He has students time their procrastination for a week and uses this as a pretest (before) indicator of procrastination. Students then attend a workshop in which they are instructed to do a specific warming-up exercise for studying by focusing on a pleasant activity. For the next week, students again time their procrastination. The counselor then uses the time from this week as the posttest (after) measure. 7. Suppose the counselor wants to examine whether there is a change of any kind (either an increase or decrease) in procrastination after attending his workshop. What would be the appropriate description of “Population 2” (the population to which the population his sample represents is being compared)? A) People whose posttest scores will be lower than their pretest scores B) People whose change scores will be greater than 0 C) People whose change scores will be 0 D) People whose change scores will be less than their pretest scores 8. Presume the counselor wants to examine whether there is a change (either an increase or decrease) in procrastination after attending his workshop. If the counselor tests 10 students using the .05 level of significance, what cutoff t score(s) will he use? (You should be able to figure this out without a table.) A) 2.62, 0, +2.62 B) +2.262 C) 2.262, 0 D) 2.262, +2.262 9. Suppose the counselor found the sum of squared deviations from the mean of the sample to be 135. Given that he tested 10 people, what would be the estimated population variance? A) 135/10 = 13.5 B) 135/9 = 15.0 C) 10/135 = .074 D) 9/135 = .067 10. A researcher conducts a study of perceptual illusions under two different lighting conditions. Twenty participants were each tested under both of the two different conditions. The experimenter reported: “The mean number of effective illusions was 6.72 under the bright conditions and 6.85 under the dimly lit conditions, a difference that was not significant, t(19) = 1.62.” Explain this result to a person who has never had a course in statistics. Be sure to use sketches of the distributions in your answer.

OpenStudy (vshiroky):

The only one I think I know is 9) b

jimthompson5910 (jim_thompson5910):

population variance = (sum of squared deviations)/N where N is the population size

OpenStudy (vshiroky):

which one are you talking about?

jimthompson5910 (jim_thompson5910):

For #9

OpenStudy (vshiroky):

so it's over 10? for some reason I thought it was N-1

jimthompson5910 (jim_thompson5910):

you're thinking of the sample variance

jimthompson5910 (jim_thompson5910):

or sample standard deviation

OpenStudy (vshiroky):

OO.. so my answer would be A then

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (vshiroky):

Gotcha.. how do I figure out the others?

jimthompson5910 (jim_thompson5910):

I'm not sure about 7. I'll have to come back to that.

jimthompson5910 (jim_thompson5910):

For #8, I think you have to find the value k that satisfies this equation P(T < k) = 0.025 with 10-1 = 9 degrees of freedom

OpenStudy (vshiroky):

What do the letters stand for?

jimthompson5910 (jim_thompson5910):

T is a random variable from the T distribution

jimthompson5910 (jim_thompson5910):

do you have a T distribution table?

OpenStudy (vshiroky):

No.. it says I should be able to figure it out without a table

jimthompson5910 (jim_thompson5910):

you either need a table or calculator

jimthompson5910 (jim_thompson5910):

there's no way to do it by hand (at least in a reasonable amount of time)

OpenStudy (vshiroky):

Think I can google a test?

jimthompson5910 (jim_thompson5910):

here's one good calculator you can use http://www.usablestats.com/calcs/tinv

OpenStudy (vshiroky):

What number do I enter for proportion? and this is 2 tailed right because they are asking for 2 numbers?

jimthompson5910 (jim_thompson5910):

you can enter 0.05 and make it 2 sided if it were one sided, then you'd enter 0.025 (since 0.05/2 = 0.025)

OpenStudy (vshiroky):

so my answer would be B

jimthompson5910 (jim_thompson5910):

keep in mind this is a two tailed test

OpenStudy (vshiroky):

Ok.. so d?

jimthompson5910 (jim_thompson5910):

good

OpenStudy (vshiroky):

Are you able to help with 7 and 10?

jimthompson5910 (jim_thompson5910):

I'm not sure about 7 still. I think the claim that the scores would be lower would indicate he wants population 2 to have lower scores, but I don't think that's how you'd define population 2

jimthompson5910 (jim_thompson5910):

I might be missing something about it

OpenStudy (vshiroky):

Alright... thank you for all your help

jimthompson5910 (jim_thompson5910):

you're welcome